# Evaluating Visual and Cultural Interpretation: The K-Viscuit Benchmark with Human-VLM Collaboration

### v1

@@ Abstract
To create culturally inclusive vision-language models (VLMs), the foremost requirement is developing a test benchmark that can diagnose the models’ ability to respond to questions reflecting cultural elements. This paper addresses the necessity for such benchmarks, noting that existing research has relied on human annotators’ manual efforts, which impedes diversity and efficiency. We propose a semi-automated pipeline for constructing cultural VLM benchmarks to enhance diversity and efficiency. This pipeline leverages human-VLM collaboration, where VLMs generate questions based on guidelines, human-annotated examples, and image-wise relevant knowledge, which are then reviewed by native speakers for quality and cultural relevance. The effectiveness of our adaptable pipeline is demonstrated through a specific application: creating a dataset tailored to Korean culture, dubbed K-Viscuit. The resulting benchmark features two types of questions: Type 1 questions measure visual recognition abilities, while Type 2 assess fine-grained visual reasoning skills. This ensures a thorough diagnosis of VLM models across various aspects. Our evaluation using K-Viscuit revealed that open-source models notably lag behind proprietary models in understanding Korean culture, highlighting areas for improvement. We provided diverse analyses of VLM performance across different cultural aspects. Besides, we explored the potential of incorporating external knowledge retrieval to enhance the generation process, suggesting future directions for improving cultural interpretation ability of VLMs. Our dataset and code will be made publicly available.
@@ 1 Introduction
The utility of vision-language models (VLMs) in bridging visual and textual data is well-recognized, offering advancements in applications such as image captioning and optical character recognition. However, most VLMs are predominantly trained on data from Western languages and cultures (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), limiting their effectiveness in non-Western contexts (Liu et al., 2021; Yin et al., 2023, 2021; Romero et al., 2024). Cultural nuances significantly influence the interpretation of visual and textual information, impacting the accuracy and relevance of VLM outputs. Consequently, building VLMs that understand diverse cultural backgrounds is essential for enhancing inclusivity and ensuring fair performance across different demographics.
To develop VLMs that effectively understand the cultural contexts of specific countries, establishing a wide-ranging test benchmark is a critical first step. Such a benchmark is essential for diagnosing the model’s ability to accurately interpret and respond to culturally specific inputs. Notably, previous research has endeavored to create VLM test benchmarks for particular countries (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024).
Figure 1: Pipeline Overview.
However, these approaches have primarily relied on manually crafted questions, which present two major limitations. First, their process is time-consuming and labor-intensive, increasing the cost of the research and making it difficult to extend and apply to the cultural contexts of other countries. Additionally, the diversity of questions can be limited due to cognitive fixation, where humans find it challenging to generate a wide range of varied questions after settling on an initial idea (Ramos, 2020). This lack of diversity can hinder the accurate diagnosis of the model’s ability to generalize across different types of culturally nuanced inputs.
To address these limitations, we propose a semi-automated pipeline for constructing cultural VLM benchmarks that enhance the efficiency and diversity of culture-relevant visual question and answer generation as illustrated in Fig. 1.
This pipeline incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers leveraging carefully crafted guidelines, a small set of human-annotated examples, and image-wise relevant knowledge. These recommendations are then assessed and approved by native speakers, ensuring high quality and cultural relevance.
In this work, we applied the proposed pipeline to create a dataset for Korean culture, named K-Viscuit (Korean Visual and Cultural Interpretation Test). Nevertheless, our pipeline can be adapted and applied to other cultures analogously. The resulting benchmark dataset created using our proposed pipeline has two notable features. Firstly, it has the ability to measure nuanced cultural understanding through two types of assessments: Type 1, which measures visual recognition capability, and Type 2, which measures visual reasoning capability. Secondly, the multiple-choice questions in the dataset include four options, with distractors generated following carefully designed instructions to ensure high similarity among the options. This high similarity between options prevents models from using shortcuts to determine the correct answer, ensuring a more robust and comprehensive evaluation of cultural understanding in VLMs. These features underscore the effectiveness and thoroughness of the resulting benchmark dataset in addressing cultural nuances.
Using our benchmark data, we evaluated the performance of various open-source and proprietary VLMs. Our experimental results indicate that open-source models significantly lag behind proprietary models in the context of Korean culture. This assessment highlights key improvement areas and sets development goals for open-source models, with proprietary models as the reference. Our contributions are summarized as follows:
  • We propose a pipeline for building benchmarks to diagnose VLMs’ understanding of various cultures. It enables efficient creation of culturally diverse datasets through human-VLM cooperation.
@@
  • Using our pipeline, we constructed the K-Viscuit Korean Culture VQA benchmark.
  • We utilized K-Viscuit to diagnose and analyze the performance of existing open-source and proprietary VLMs.
@@ 2 Related Work
Recent research has made significant strides in developing benchmarks to assess the ability of AI models, particularly VLMs and Large Language Models (LLMs), to understand and respond appropriately to various cultural contexts. These efforts are crucial given that many existing models are trained predominantly on Western-centric datasets (Young et al., 2014; Lin et al., 2014; Antol et al., 2015), which constrains their performance in non-Western settings. On the side of LLMs, Kim et al. (2024) introduced a benchmark dataset, CLIcK, specifically for assessing cultural and linguistic intelligence in Korean language models. This dataset includes QA pairs designed to test the cultural knowledge of Korean language models. Wibowo et al. (2023) created COPAL-ID, a dataset that captures Indonesian cultural nuances for text-based commonsense reasoning, highlighting the need for culturally aware benchmarks.
VLMs require consideration of both image and text inputs to effectively reflect cultural contexts, as cultural nuances can arise from either source. Liu et al. (2021) proposed the MaRVL dataset, a multilingual visually grounded reasoning dataset that includes five different languages and cultures to diagnose inherent language and cultural biases in pre-trained models. Their work highlights the need for models to seamlessly interpret and integrate visual and textual information across diverse cultures, paving the way for more inclusive and globally aware AI systems. Yin et al. (2021) built the GD-VCR dataset, which tests visual commonsense reasoning influenced by geographical and cultural contexts. Romero et al. (2024) presented the CVQA, a multilingual visual question-answering (VQA) benchmark that incorporates various languages and cultures, providing a comprehensive tool for evaluating VLMs in diverse cultural contexts.
Previous efforts to create cultural benchmarks for VLMs, however, have heavily relied on human annotators, which can limit both the diversity and efficiency of dataset creation (Liu et al., 2021; Yin et al., 2021; Ramaswamy et al., 2024; Romero et al., 2024). The efficiency and diversity of dataset creation can be significantly enhanced by leveraging AI models, particularly when combined with human collaboration. This approach not only improves the speed and scope of dataset generation but also maintains high-quality standards. Liu et al. (2022) demonstrated this in their study on building a natural language inference dataset, where they combined the generative capabilities of language models with human review. Analogous to how LLMs are heavily utilized in the creation of datasets for language-only tasks (Taori et al., 2023; Kim et al., 2023; Dubois et al., 2024; Kim et al., 2024), leveraging the inherent knowledge of AI models in building multimodal cultural benchmarks can yield diverse and culturally nuanced examples. In this work, by harnessing AI models, we have significantly enhanced the diversity and richness of the dataset, ensuring it better reflects the wide array of cultural contexts we aim to represent.
3 Data Construction Pipeline
@@
We present our human-AI collaborated dataset construction pipeline to diagnose the understanding of VLMs in a target cultural field. In this work, we focus on Korean culture as a target cultural field. First, we provide an overview (section 3.1) and details of the pipeline. Then, we present the analysis of the resulting dataset, dubbed K-Viscuit (Korean Visual and Cultural Interpretation Test) (section 3.2).
3.1 Pipeline Overview
@@
The pipeline is designed to comprise a multiple-choice visual question answering (VQA) task, where each evaluation sample consists of an image, a question, and four options with one correct choice. Native Korean speakers are involved in the dataset construction, and a powerful proprietary VLM is employed to mitigate unintended human biases, such as cognitive fixation (Ramos, 2020), and encourage an effective annotation process. The generated samples are designed to cover various aspects of a target culture derived from daily life and necessitate multimodal reasoning to interpret both the image and the question accurately. All text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects. The dataset construction consists of four stages: 1) concept selection, 2) image selection, 3) question and options annotation, and 4) human verification. Fig. 1 illustrates the overall pipeline. The details in each stage are as follows.
@@ 3.1.1 Concept Categorization
Inspired by recent studies on constructing multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge related to various concepts that Korean natives encounter in daily life. Each concept should not be exclusive to a specific culture but should have some degree of universality. To define the list of such concepts, we refer to semantic concepts provided in the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015), following Liu et al. (2021). Based on the concepts in the resource and discussions, our dataset is based on top of the following ten concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
@@ 3.1.2 Image Selection
Based on the list of chosen concepts above, Korean native annotators manually collected images from the web. To avoid an excessive number of images containing specific objects, each object is limited to appearing a maximum of two times within any single category. This encourages a broad representation of concepts related to Korean culture. Moreover, following Liu et al. (2021), only images depicting concepts that could physically exist in everyday life were selected. Annotators are encouraged to creatively choose diverse and suitable images by using various web resources. Wikimedia Commons [^1] https://commons.wikimedia.org/wiki are used as the primary sources, and only CC-licensed images are selected.
@@ 3.1.3 Question Generation
Based on the selected images, we annotate questions in the form of multiple-choice QA. To comprehensively evaluate the understanding of Korean culture, we categorize the questions into two types: visual recognition (Type 1) and reasoning (Type 2). The visual recognition questions ask about basic visual information (e.g., counting, color recognition) or the names of objects appearing in the image. The reasoning questions go beyond simple visual recognition and ask for fine-grained knowledge or reasoning processes related to the given image. For each image, a single Type 1 question and between one and four Type 2 questions were created. This categorization offers two main advantages. First, the Type 1 questions allow for a multifaceted assessment of how well a model understands an image and its embedded concepts. Second, the Type 2 questions ensure that our dataset comprehensively evaluates the understanding of various aspects of Korean culture, rather than merely falling back to simple object recognition.
@@
AI-assisted Question Annotation
We create questions and their options (including one correct answer and three distractors) by leveraging a powerful proprietary large VLM (i.e., GPT-4-Turbo). For each concept category, human annotators first write questions and options for at least three images in each category. These manually annotated examples serve as demonstrations for the VLM to generate questions and options along with additional information.
In detail, the VLM receives 1) the target image, 2) a set of human-annotated demonstration examples for other images, 3) a detailed guideline for annotation, and 4) an image-wise textual knowledge description. We include a description or general knowledge associated with the given image to enhance the diversity and relevance of the questions. This ensured that the VLM could generate questions grounded in real-world knowledge. One important note is that the guidelines emphasize promoting similarity among all four options for multiple-choice questions, a principle kept in mind while creating the human-annotated examples, inspired by Wang et al. (2023). All textual information is delivered to the VLM through a natural language prompt, and different annotation guidelines are used for generating questions in different types (i.e., visual recognition and reasoning). The text prompts for VLM are presented in Appendix A.
@@ 3.1.4 Human Verification
During our preliminary study, we observed that the model often generated inappropriate data samples. For instance, some questions are open-ended or do not induce a single clear answer. In other cases, generated questions are subjective or not sufficiently related to the given image or Korean culture. To ensure the quality of automatically generated datasets, human verification is performed on the generated samples from the VLM. The VLM is guided to generate five question-option sets per image. If none of the five generated samples for a single image meet the criteria, we ask the VLM to regenerate questions and answers. The generated samples deemed appropriate are included in the final dataset with minimal revisions if required. We note that this AI-assisted annotation approach offered several advantages: It significantly accelerated the annotation process and helped minimize unintended human biases. By including human verification on generated questions, we ensure both accuracy and cultural relevance.
@@ 3.2 K-Viscuit
The detailed statistics of our final benchmark dataset are shown in Table 1. Our dataset comprises a total of 657 examples, with 237 Type 1 questions and 420 Type 2 questions, drawn from 237 unique images from 10 concept categories. The average word count for questions in Type 1 and Type 2 are 10.11 and 15.46, respectively. The average word count for all questions is 13.53 words. Each question is accompanied by four options, resulting in a total of 2,628 options with an average word count of 1.74 words. Fig. 2 illustrates the distribution of selected concepts of samples in our dataset.
@@
To better understand the characteristics of our dataset, we categorize the types of required knowledge to solve questions. We analyzed Type 2 questions using the GPT-4 model following Tong et al. (2024). All Type 2 QA pairs were provided, and the knowledge required to answer each question was summarized. As shown in Fig. 4, diverse cultural elements were considered. Detailed instructions for the categorization can be found in the Appendix B.1.
Figure 4: Required Cultural Knowledge.
@@
Fig. 3 presents sampled images, questions, and options along with their concept category and question type.

### v2

@@ Abstract
To create culturally inclusive vision-language models (VLMs), developing a benchmark that tests their ability to address culturally relevant questions is essential. Existing approaches typically rely on human annotators, making the process labor-intensive and creating a cognitive burden in generating diverse questions. To address this, we propose a semi-automated framework for constructing cultural VLM benchmarks, specifically targeting multiple-choice QA. This framework combines human-VLM collaboration, where VLMs generate questions based on guidelines, a small set of annotated examples, and relevant knowledge, followed by a verification process by native speakers. We demonstrate the effectiveness of this framework through the creation of K-Viscuit, a dataset focused on Korean culture. Our experiments on this dataset reveal that open-source models lag behind proprietary ones in understanding Korean culture, highlighting key areas for improvement. We also present a series of further analyses, including human evaluation, augmenting VLMs with external knowledge, and the evaluation beyond multiple-choice QA. [^1] Our dataset is available at https://huggingface.co/datasets/ddehun/k-viscuit.
@@ 1 Introduction
Recent advances in vision-language models (VLMs) have demonstrated remarkable capabilities in tasks ranging from image captioning to visual question answering. However, these models are predominantly trained on Western-centric datasets (Lin et al., 2014; Young et al., 2014; Antol et al., 2015), leading to significant performance disparities when applied to non-Western contexts (Liu et al., 2021; Yin et al., 2021, 2023; Romero et al., 2024). This cultural bias is particularly problematic as visual interpretation often depends heavily on cultural context, necessitating the development of more culturally aware VLMs.
Several benchmarks have been proposed to evaluate cultural understanding in VLMs (Liu et al., 2021; Yin et al., 2021; Romero et al., 2024; Nayak et al., 2024). These approaches primarily rely on manual question generation, which, while valuable, faces certain practical challenges. The manual process can be time-consuming and resource-intensive when scaling to new cultural contexts. Additionally, as noted in cognitive science research (Ramos, 2020), human annotators may experience cognitive fixation, potentially limiting the diversity of generated questions. These practical considerations motivate the need for more efficient benchmark construction approaches.
Figure 1: Framework Overview.
Inspired by recent successes in human-LLM collaborative data generation (Liu et al., 2022; Bartolo et al., 2022; Kamalloo et al., 2023), we propose a semi-automated framework for constructing cultural VLM benchmarks that enhances both the efficiency and diversity of culture-relevant visual question and answer generation, as shown in Fig. 1. Our framework incorporates human-VLM collaboration, where the VLM generates and recommends questions and answers based on carefully crafted guidelines, a small set of human-annotated examples, and image-specific knowledge. Native speakers then verify these recommended questions to ensure quality and cultural relevance.
Using this framework, we develop K-Viscuit (Korean Visual and Cultural Interpretation Test), a benchmark dataset for Korean culture that can be adapted for other cultural contexts. K-Viscuit features two distinct evaluation types: visual recognition and visual reasoning. Also, the benchmark employs carefully designed multiple-choice questions with highly similar distractors to prevent models from exploiting superficial patterns.
Our evaluation with K-Viscuit reveals a significant performance gap between open-source and proprietary VLMs in understanding Korean culture. We provide insights into the current limitations and potential improvements in VLMs’ cultural understanding capabilities through detailed analyses, which include human evaluation, external knowledge integration, and extended evaluation beyond a multi-choice question-answering setup.
Our contributions are summarized as follows:
  • We propose a semi-automated framework for constructing benchmarks to evaluate the cultural understanding capabilities of VLMs.
@@
  • We develop K-Viscuit, a Korean culture-focused VQA benchmark using our proposed framework.
  • We present comprehensive experimental results and analyses of both open-source and proprietary VLMs evaluated on K-Viscuit.
@@ 2 Related Work
Recent research has made significant strides in developing benchmarks to assess AI models’ cultural understanding capabilities. These efforts are particularly important as many existing models, including VLMs and Large Language Models (LLMs), are trained predominantly on Western-centric datasets (Young et al., 2014; Lin et al., 2014; Antol et al., 2015), limiting their effectiveness in non-Western contexts. For LLMs, several notable cultural benchmarks have emerged: Kim et al. (2024) introduced CLIcK, a benchmark for evaluating Korean language models’ cultural knowledge through carefully designed QA pairs, while Wibowo et al. (2023) developed COPAL-ID to capture Indonesian cultural nuances in text-based commonsense reasoning.
In the multimodal domain, VLMs require consideration of both visual and textual inputs to effectively reflect cultural contexts. Liu et al. (2021) addressed this challenge with MaRVL, a multilingual visually-grounded reasoning dataset spanning five languages and cultures, demonstrating the importance of cross-cultural visual-linguistic understanding. Building on this foundation, Yin et al. (2021) introduced GD-VCR to evaluate geographical and cultural aspects of visual commonsense reasoning, while Romero et al. (2024) developed CVQA, a comprehensive multilingual VQA benchmark for assessing VLMs across diverse cultural contexts.
While these cultural benchmarks have provided valuable insights, their reliance on manual annotation can constrain both the diversity and efficiency of dataset creation (Liu et al., 2021; Yin et al., 2021; Ramaswamy et al., 2024; Romero et al., 2024). Recent work has demonstrated the potential of AI-assisted dataset generation when combined with human expertise. Liu et al. (2022) successfully applied this approach to natural language inference tasks, and similar strategies have been widely adopted in creating language-only datasets (Taori et al., 2023; Kim et al., 2023; Dubois et al., 2024; Kim et al., 2024). Our work extends this paradigm to multimodal cultural benchmarks, leveraging AI models to enhance dataset diversity while maintaining high quality through human verification.
3 Data Construction Framework
@@
We present our human-AI collaborative framework for constructing datasets to evaluate VLMs’ understanding of specific cultural domains. In this work, we focus on Korean culture as our target domain. First, we provide an overview (section 3.1) and implementation details. Then, we present the analysis of our resulting dataset, K-Viscuit (Korean Visual and Cultural Interpretation Test) (section 3.2).
3.1 Framework Overview
@@
Our framework is designed to create a multiple-choice visual question answering (VQA) task, where each evaluation sample consists of an image, a question, and four options with one correct answer. Native Korean speakers participate in the dataset construction process, while a powerful proprietary VLM is employed to mitigate unintended human biases, such as cognitive fixation (Ramos, 2020), and streamline the annotation process. The generated samples cover various aspects of the target culture derived from daily life and require multimodal reasoning to interpret both visual and textual information accurately. The dataset construction consists of four stages: 1) concept selection, 2) image selection, 3) question and options annotation, and 4) human verification. Fig. 1 illustrates the overall framework.
@@ 3.1.1 Concept Categorization
Inspired by recent studies on multicultural evaluation datasets (Liu et al., 2021; Wibowo et al., 2023; Kim et al., 2024), we aim to assess knowledge of various concepts encountered in daily life by Korean natives. While each concept should have some degree of universality, its manifestation often varies across cultures. Following Liu et al. (2021), we reference semantic concepts from the Intercontinental Dictionary Series (IDS) (Key and Comrie, 2015) to define our concept list. Our dataset encompasses ten core concepts: Food, Beverage, Game, Celebrations, Religion, Tool, Clothes, Heritage, Architecture, and Agriculture.
@@ 3.1.2 Image Selection
Korean native annotators collected web images corresponding to the selected concepts. To ensure diverse representation, we limited each specific object to appearing no more than twice within any single category. Following Liu et al. (2021), we selected only images depicting concepts that could physically exist in everyday life. Annotators were encouraged to source diverse and suitable images from various web resources. Wikimedia Commons [^2] https://commons.wikimedia.org/wiki served as the primary source, and only CC-licensed images were selected.
@@ 3.1.3 Question Generation
Based on the selected images, we annotate questions in multiple-choice QA format. To comprehensively evaluate understanding of Korean culture, we categorize questions into two types: visual recognition (Type 1) and reasoning (Type 2). Visual recognition questions assess basic visual information such as object identification, while reasoning questions require fine-grained cultural knowledge or deeper reasoning processes related to the image. For each image, we created one Type 1 question and between one to four Type 2 questions. This categorization offers two key advantages: First, Type 1 questions enable assessment of a model’s basic visual understanding of culturally embedded concepts. Second, Type 2 questions comprehensively evaluate cultural understanding beyond simple object recognition.
@@
AI-assisted Question Annotation
We create questions and their options (one correct answer and three distractors) by leveraging a powerful proprietary VLM (GPT-4-Turbo). For each concept category, human annotators first create exemplar questions and options for at least three images. These manually annotated examples serve as demonstrations for the VLM to generate additional questions and options.
Specifically, the VLM receives: 1) the target image, 2) human-annotated demonstration examples, 3) detailed annotation guidelines, and 4) image-specific knowledge descriptions. We include relevant contextual knowledge for each image to enhance question diversity and relevance, ensuring VLM-generated questions are grounded in real-world understanding. Notably, following Wang et al. (2023), our guidelines emphasize maintaining high similarity among all four multiple-choice options, a principle also reflected in human-annotated examples. All information is provided to the VLM through natural language prompts, with distinct annotation guidelines for visual recognition and reasoning questions. The detailed prompts are presented in Appendix A.
It should be noted that all text in the dataset is written in English to isolate the evaluation of multicultural comprehension from multilingual aspects. However, as culture and language are not entirely orthogonal (SUSAN, 1996; Kramsch, 2014), completely separating them can be challenging. For instance, we observed that certain Korean terms lack exact English equivalents. In such cases, we romanize these Korean terms following standard transliteration rules.
@@
Model All Food Beverage Game Celeb. Religion Tool Clothes Heritage Arch. Agri. InstructBLIP-7B (Dai et al., 2024) 50.84 40.85 42.31 38.46 53.19 40.74 50.67 62.16 51.61 60.55 72.22 instructBLIP-13B (Dai et al., 2024) 55.56 45.77 50.00 46.15 59.57 55.56 54.67 64.86 66.13 60.55 64.81 mPLUG-Owl2-7B (Ye et al., 2023) 48.25 42.25 42.31 30.77 63.83 55.56 48.00 54.05 45.16 49.54 66.67 LLaVA-1.6-7B (Liu et al., 2024) 56.32 43.66 48.08 40.38 57.45 51.85 54.67 67.57 59.68 72.48 72.22 LLaVA-1.6-13B (Liu et al., 2024) 57.08 45.07 53.85 36.54 68.09 40.74 53.33 70.27 66.13 69.72 70.37 InternLM-XC2-7B (Dong et al., 2024) 59.67 50.70 48.08 40.38 65.96 55.56 58.67 64.86 69.35 69.72 75.93 Molmo-7B-D (Deitke et al., 2024) 61.04 58.45 71.15 44.23 68.09 44.44 61.33 70.27 56.45 64.22 68.52 Idefics2-8B (Laurençon et al., 2024) 63.62 51.41 50.00 50.00 74.47 66.67 69.33 75.68 74.19 73.39 62.96 Llama-3.2-11B (Dubey et al., 2024) 68.04 61.27 65.38 50.00 72.34 70.37 72.00 75.68 72.58 69.72 81.48 Claude-3-opus (Anthropic, 2024) 70.02 62.68 73.08 59.62 72.34 77.78 74.67 78.38 75.81 67.89 75.93 GPT-4-Turbo (Achiam et al., 2023) 80.82 73.94 80.77 78.85 85.11 85.19 81.33 86.49 85.48 79.82 87.04 Gemini-1.5-Pro (Reid et al., 2024) 81.58 80.28 78.85 71.15 85.11 77.78 82.67 83.78 83.87 84.40 85.19 GPT-4o (OpenAI, 2024) 89.50 88.73 82.69 86.54 95.74 85.19 90.67 91.89 91.94 91.74 87.04
@@ 3.1.4 Human Verification
Our preliminary studies showed that while the VLM often produced factually correct samples, some did not fully align with the cultural nuances we sought to capture. Rather than discarding only incorrect content, our human verification process prioritizes selecting question-option sets that best reflect intended cultural subtleties. Although this leads to setting aside numerous plausible samples, it does not imply unreliability; instead, it refines the dataset to ensure deep cultural resonance. Approved samples require minimal revision before inclusion. This approach emphasizes nuanced cultural alignment and indicates the need for further work to better synchronize VLM outputs with human cultural intentions.
Figure 4: Required Cultural Knowledge.
@@ 3.2 K-Viscuit
Table 1 presents detailed statistics of our benchmark dataset. Our dataset comprises 657 total examples (237 Type 1 and 420 Type 2 questions) based on 237 unique images across 10 concept categories. The average word counts are 10.11 and 15.46 for Type 1 and Type 2 questions respectively, with an overall average of 13.53 words. Each question includes four options, totaling 2,628 options with an average length of 1.74 words. Fig. 3 shows the distribution of concept categories in our dataset. We also compare our dataset size with CVQA (Romero et al., 2024), a recently proposed cultural VQA benchmark, in Appendix E.
@@
To characterize our dataset, we analyze the types of knowledge required to solve questions. Following Tong et al. (2024), we used GPT-4 to analyze Type 2 questions. We provided all Type 2 QA pairs and summarized the required knowledge for each question. As shown in Fig. 4, the analysis reveals diverse cultural elements. Detailed categorization instructions are provided in Appendix B.1.
@@
Fig. 2 showcases sample images, questions, and options along with their concept categories and question types.